All Terrain Tires manufactures three different off-road ATV tires: Model A, Model B, and Model C. Plenty of market demand exists for all models. The table below reports the prices and costs per unit of each product:

Model A

Model B

Model C

Selling price

$ ?

$50

$55

Direct materials costs

$5

$5

$5

Direct labor costs ($15 per labor hour)

$15

$15

$30

Variable support costs ($3 per machine hour)

$3

$6

$9

Fixed support costs

$10

$10

$10

Assuming that machine hours are limited (i.e., this is the constrained resource), at what price does Model A become the most profitable model to produce?

Group of answer choices

A. Any price above $55

B. Any price above $47

C. Any price above $23

D. Any price above $35
Answer #1

Contribution for MODEL B = Selling price - direct materials cost - direvt labor costs - variable support costs = $50 - $5 - $15 - $6 = $24.

Contribution per machine per hour = contribution / machine hours = $24 / 2 hours = $8 per hour.

Contribution for model c = $55 - $5 - $30 - $9 = $11

Contribution per machine hour = $11 / 3 hours = $3.67 per hour.

Variable costs for model A = $5 + $15 + $3 = $23 .

Contribution for model A will be maximum when the selling price is maximum.

Therefore, selling price of $55 will be most profitable.

Contribution per machine hour at $55 selling price = ( $55 - $23 ) / 1 hour = $32 per hour.

Therefore, at this selling price Model A Will be most profitable.

Machine hours are calculated by formula = variable support costs / $3 per machine hours.